OpenStudy (anonymous):

Which formula do you use for this one? @jdoe0001

OpenStudy (jdoe0001):

hmm have you covered logarithms yet?

OpenStudy (anonymous):

No.

OpenStudy (jdoe0001):

well... this one involves compound interest and you'd need to find how long is "t" or years in \(\bf A=P\left(1+\frac{r}{n}\right)^{nt}\) which means you'd end up with a logarithmic equation so if you haven't covered that yet... you may want to cover it firstly

OpenStudy (anonymous):

Oh I thought you meant \[\log_{2} \] that stuff

OpenStudy (anonymous):

@jdoe0001 What is n?

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\\ r=rate\to r\%\to \frac{r}{100}\\ n=\textit{times it compounds per year}\\ t=years\)

OpenStudy (jdoe0001):

so if your interest rate compounds semi-annually, that means twice a year thus n = 2 if it compounds monthy, that means 12 times in 1 year, thus n = 12 if it compounds bimonthly, that means 6 times in 1 year so n = 6 if it compounds every two years, that means in 1 year is 1/2 n = 0.5 and so on

OpenStudy (anonymous):

6,449.82\[(1+\frac{ 6.9 }{ 0.5 }) ^.5t\]

OpenStudy (jdoe0001):

well. yes... but "t" is an exponent... that means you'd need to use a logarithm function

OpenStudy (anonymous):

don't you just solve for "t" ?

OpenStudy (jdoe0001):

yes you'd, after taking logarithm to both sides, with certain base but you haven't covered that

OpenStudy (anonymous):

6449.82(14.8) ^0.5t

OpenStudy (jdoe0001):

well.... if you havent' covered logarithms.... I'd think this one wouldn't apply to yyoyu

OpenStudy (anonymous):

I have to do it, it does apply.

OpenStudy (anonymous):

May you walk me through it?

OpenStudy (jdoe0001):

well... I have to dash for one.... but logarithms can be a whole chapter, or at least a few sections so... you can find good material though https://www.khanacademy.org/math/algebra2/logarithms-tutorial

OpenStudy (anonymous):

:/